OpenStudy (anonymous):

find the derivative of the function. g(t)=e (to the power of-3/t^2)

OpenStudy (anonymous):

\[e^{\frac{-3}{t^2}}\]?

OpenStudy (anonymous):

is it -3e to the power of t^2?

OpenStudy (anonymous):

since the derivative of \(e^x\) is \(e^x\) your job is only to find the derivative of \(-\frac{3}{t^2}\) and multiply it by \[e^{-\frac{3}{t^2}}\]

OpenStudy (anonymous):

ok so i find what -3/t^2 is and then multiply it by e^x right?

OpenStudy (anonymous):

no not by \(e^x\) but rather by \(e^{-\frac{3}{t^2}}\)

OpenStudy (anonymous):

don't change the input for the "outside" function, leave it alone

OpenStudy (anonymous):

o okay so it would be this -3/t^2 =====-3(t^-2) and then i could use the power rule right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

kkkkkkkkk! so i would get 0 times t^-2 +-2t (0) right?

OpenStudy (anonymous):

i mean -2t times -3

OpenStudy (anonymous):

right?